# On nonlinear transitions, minimal seeds and exact solutions for the geodynamo

### v1

@@ 3.1 Benchmark setup
For the purpose of the present study, and to test the numerical scheme, we adopt the parameters of the dynamo benchmark introduced by Christensen et al. (2001): $\textit{Ek}=10^{-3}$, $\tilde{\textit{Ra}}=100$, $\textit{Pm}=5$, $\textit{Pr}=1$ and $r_{i}/r_{o}=0.35$. This benchmark is a widely studied subcritical dynamo solution, and therefore provides an ideal case on which to test our optimisation procedure. The initial temperature field for the benchmark is specified as
@@
with $x=2r-r_{i}-r_{o}$, and $a=0.1$, driving an $m=4$ convective instability (‘case 0’ of Christensen et al. (2001): non-magnetic convection). While the resulting hydrodynamic state is linearly stable to magnetic perturbations, it is (nonlinearly) unstable to some finite-amplitude magnetic perturbations that trigger dynamo action. Christensen et al. (2001) suggest that the basin of attraction of this dynamo solution may be small and therefore recommend prescribing a mainly dipolar, initial magnetic field of the form
@@
with (dimensionless) magnetic energy $M_{0}=1215$, which kickstarts an $m=4$ dynamo solution. We note that, as at the benchmark parameters the system is hydrodynamically unstable to multiple modes of convection, other dynamo solutions are possible using different thermal forcing wavenumbers — see Feudel et al. (2017) for instance.
K M $\omega$ Stable
$58.35$ 00 $0.000182$ Yes
$30.773$ $626.4$ $-0.00310$ Yes
$29.72$ $459.9$ $-0.00250$ No
Table 1: Travelling wave solutions to the governing equations with $m=4$ symmetry. $\omega$ is the frequency at which they drift relative to the reference frame.
@@
Figure 2: Slices of the $m=4$ nonlinear dynamo state (TW1).
@@
Using the Newton-hookstep procedure outlined in §2.3, we find the system to possess three $m=4$ travelling wave (TW) solutions at our working parameters, characterised by the energies and drift frequency (relative to the already-rotating reference frame) displayed in table 1. The table shows one purely hydrodynamic solution (TW0) which corresponds to the ‘case 0’ benchmark of Christensen et al. (2001). This hydrodynamic solution arises via a supercritical Hopf bifurcation at $\tilde{\textit{Ra}}_{c}=55.9$, and is linearly stable to both hydrodynamic and magnetic perturbations at our working parameters. The two other TW solutions are MHD solutions that form at a saddle node bifurcation at $\tilde{\textit{Ra}}\sim 98$ (Feudel et al., 2017) connected to the purely convecting solution: one stable (TW1) and corresponding to the dynamo (‘case 1’ of Christensen et al. (2001)); the other unstable (TW2). Figure 2 shows slices of TW1, showing that the nonlinear state is mainly dipolar, with a quadrupolar toroidal part (evident from the radial component of the current). We note here that TW2 has a very similar structure to TW1, which is to be expected since we are close to the saddle node bifurcation, although it shows slightly more localisation than TW1. This localisation becomes more apparent at higher Rayleigh numbers, where the two states start to deviate in structure.
@@
We now let the system evolve to the saturated, purely hydrodynamic, $m=4$ convective solution (TW0). This state is then seeded with a magnetic field of the form given by Christensen et al. (2001) (31)-(32); however, the field amplitude is rescaled to have a given initial magnetic energy $M_{0}$, which is gradually decreased (through a series of repeated simulations) down to the point where the initial field fails to trigger a dynamo. The energy timeseries for two slightly different energy budgets close to the threshold are shown in figure 3. Initially, the evolution is very similar for both cases and is characterised by a transient amplification of the magnetic field, which extracts energy from the convection. This transient growth occurs on a dynamical time scale and is followed by a plateau until around $t^{*}\approx 2.5$ (i.e. half an ohmic timescale). At this point the evolutions differ, with the $M_{0}=345$ case achieving further growth in magnetic energy to reach a new stable MHD state (corresponding to the TW1 dynamo) over a few ohmic timescales, and the $M_{0}=344$ case magnetic field decaying back to the purely hydrodynamic state TW0. Remarkably, the flow state around $t^{*}$ corresponds to the unstable MHD state TW2: it is thus found to be an unstable attractor, with both solutions initially approaching TW2 along its stable directions. Lastly, TW2 clearly has multiple unstable directions, with one leading to the growth of the magnetic field (and eventually the dynamo TW1), and the other leading to the recovering of the kinetic energy and the decay of the magnetic field.
@@ 3.2 Optimised seeds
With the baseline behaviour for our system established, we now seek to identify the spatial structure of the lowest-energy magnetic perturbation required to reach the dynamo solution, using the optimisation procedure outlined in §$\S$2.2. Following Mannix et al. (2022), we initially consider to that end the cost functional
$\mathcal{J}_{I}=\frac{1}{2\textit{Ek}\textit{Pm}}\langle\bm{B},\bm{B}\rangle,$ (33)
@@
which corresponds to the time-integrated magnetic energy. As our expected magnetic field is large scale, we initialise the optimisation procedure with a random guess with spherical harmonic degrees up to $\ell=m=2$, and a random radial dependence with Jacobi polynomial degree up to four. We note here, that we have tested random initial guesses up to $\ell=m=4$. The behaviour of the optimisation is initially similar, with the seed magnetic field converging to a large scale solution. However, the less sparse spectra of the $\ell=m=4$ guess causes slower convergence, with the optimisation spending a long time removing finer scale components from the initial guess. This problem is common, and could be exacerbated by the fact that our continuous gradient is not exactly the true gradient of the numerical quantity we calculate (see Mannix et al. (2024), for example). We further describe the convergence behaviour of the optimisation routine in appendix C. Based on the quick evolution of transients in the previous section, we start by considering a short optimisation window of $T_{\textit{opt}}=0.2$. The initial magnetic energy budget is specified to be $M_{0}=344$ as this is below the threshold at which a dynamo can be reached from the baseline solution.
@@
Figure 4: Comparison of the magnetic energy (top) and kinetic energy (bottom) evolution with four different initial conditions with $M_{0}=344$. The initial conditions are the optimised seeds obtained by optimising the total magnetic energy with a random/ rescaled benchmark guess (solid blue line/ dotted red line), rescaled benchmark (dashed orange line), and optimised seed obtained by optimising the energy in the $m=0$ part of the magnetic field starting from a random guess (dashed-dotted green line). The plots on the left show the energy evolution in the optimisation window $T_{\textrm{opt}}=0.2$, and the plots on the right show the long-time evolution.
(a) Meridional slice of the radial component of the magnetic field.
(b) Meridional slice of the radial component of the current.
(d) Meridional slice of the radial component of the magnetic field.
(e) Meridional slice of the radial component of the current.
(f) Equatorial slice of the latitudinal component of the magnetic field.
(g) Meridional slice of the radial component of the magnetic field.
(h) Meridional slice of the radial component of the current.
(i) Equatorial slice of the latitudinal component of the magnetic field.
@@
Figure 5: Top: Slices of an example of optimal seed obtained with the total magnetic energy cost functional, with $T_{\textrm{opt}}=0.2$ and $\textit{M}_{0}=344$. Middle: Slices of the optimal seed obtained with the axisymmetric magnetic energy cost functional with $T_{\textrm{opt}}=0.2$ and $\textit{M}_{0}=344$. Bottom: Slices of the optimal seed obtained with the axisymmetric magnetic energy cost functional with $T_{\textrm{opt}}=0.4$ and $\textit{M}_{0}=162$.
@@
Increasing the time horizon of the optimisation procedure up to $T_{\textrm{opt}}=0.4$ was not found to add robustness to the computation of optimal seeds as multiple local extrema characterised by strong transient growth (but not necessarily a long-time dynamo) continue to coexist at such target times. While it is tempting to further increase $T_{\textrm{opt}}$ — indeed we expect that optimising over (very) long times would ultimately discriminate between dynamo seeds and magnetic perturbations yielding only transient growth and ultimately decay -, for long target times convergence becomes increasingly difficult (numerical cost notwithstanding) due to the accumulation of errors in computing the gradient. It is therefore more realistic to define another cost functional that could possibly reduce the number of local extrema, yielding more robust seed identification.
@@
$\mathcal{J}_{I}=\frac{1}{2\textit{Ek}\textit{Pm}}\langle\bar{\bm{B}},\bar{\bm{B}}\rangle,$ (34)
@@
The result of optimising this cost functional is also shown in figure 4 (dashed-dotted, green line). In contrast to optimising the total magnetic energy, optimising this new cost functional is robust in the sense that it was found to always identify the same optimal seed (regardless of the random initial guess used to initialise the algorithm). Furthermore, the computed optimal seed does indeed lead to a dynamo. Incidentally, it is also found to outperform some “optimal” seeds that (locally) maximise the total magnetic energy (as exemplified in figure 4). Based on these observations, one could perhaps speculate that using this new cost functional might be a way to ‘smooth out’ the optimisation landscape, possibly suppressing the multiple local extrema and thus providing a good proxy for the global extremum of the first (total energy) optimisation problem — although this is highly uncertain.
@@
Figure 6: Robustness of the optimisation results with respect to the time horizon $T_{\textrm{opt}}$. Left: The minimum initial magnetic energy budget to find a dynamo solution for a given time horizon $T_{\textrm{opt}}$ ($m=0$ cost functional). Circles show runs that succeed in finding a dynamo, and crosses indicate no dynamo was found. Right: Magnetic energy timeseries for $M_{0}=165$ starting from optimised seeds ($m=0$ cost functional) computed with different time horizons (time horizons displayed as shaded areas)
@@
With a robust cost functional identified, we now turn our attention to lowering the energy of the initial seed. To this end, we perform a series of optimisations maximising the total axisymmetric energy, for gradually decreasing initial energy budgets at fixed time horizon. Each optimal seed then serves as initial condition for a long-time DNS: if a dynamo state is found at long times, then $M_{0}$ is decreased further and a new optimisation is performed. We thus determine the initial energy $M_{0}^{*}$ below which the computed optimal seed no longer triggers a subcritical dynamo, and repeat the whole procedure for increasing time horizon. As shown in Figure 6 (Left), $M_{0}^{*}$ was found not to vary much as the target time was increased up to $T_{\textrm{opt}}=0.4$. Moreover, the effect of lengthening the optimisation time horizon is shown in Figure 6 (Right) for an initial energy budget of $\textit{M}_{0}=165$: at this energy $\textit{T}_{\textrm{opt}}=0.2$ is not long enough to optimise an initial condition which yields a dynamo. We see that lengthening the time horizon to $\textit{T}_{\textrm{opt}}=0.3$ and $\textit{T}_{\textrm{opt}}=0.4$ enables the optimisation procedure to converge to a dynamo solution. Although the early timeseries in Figure 6 (Right) show that there are differences in the magnetic energy evolution as the $\textit{T}_{\textrm{opt}}$ is changed, we see that these changes become significantly less as $\textit{T}_{\textrm{opt}}$ is increased. This indicates that the computed extremum eventually becomes insensitive to $\textit{T}_{\textrm{opt}}$, which we can attribute to the fact that for $\textit{T}_{\textrm{opt}}\approx 0.4$ the transient growth regime has ended and the flow has approached the unstable MHD state TW2. For longer time-horizons the optimisation procedure struggles to converge.
@@
Based on these findings we conclude that $\textit{T}_{\textrm{opt}}=0.4$ is the most practical choice of time horizon, enabling the optimisation procedure to converge, while at the same time being long enough to ensure robustness with respect to changes in $\textit{T}_{\textrm{opt}}$. Recall, this is non-dimensionalised with respect to the viscous diffusion time, and this is a dynamical timescale. In other words, we only need to optimise up to when the state approaches the unstable edge state, at which point we require the instability of the edge state to lead to the dynamo solution. This is similar to previous studies finding minimal seeds for flows with edge states (see the work of Juniper (2011) for example). The spatial structure of the ‘minimal’ seed (i.e. the lowest-energy dynamo seed we could find) is shown in figure 5 (Bottom) ; it is similar to the optimal seed found at higher energy $\textit{M}_{0}=344$ (figure 5, Middle), however shows more localisation. We note that neither of these magnetic fields resemble either the nonlinear dynamo state (figure 3) or unstable state (similar to figure 2), showcasing the need for a systematic procedure that can work from random initial guesses (Mannix et al., 2022).

### v2

@@ 3.1 Benchmark setup
For the purpose of the present study, and to test the numerical scheme, we adopt the parameters of the dynamo benchmark introduced by Christensen et al., (2001): $\textit{Ek}=10^{-3}$, $\tilde{\textit{Ra}}=100$, $\textit{Pm}=5$, $\textit{Pr}=1$ and $r_{i}/r_{o}=0.35$. This benchmark is a widely studied subcritical dynamo solution, and therefore provides an ideal case on which to test our optimisation procedure. The initial temperature field for the benchmark is specified as
@@
with $x=2r-r_{i}-r_{o}$, and $a=0.1$, driving an $m=4$ convective instability (‘case 0’ of Christensen et al., (2001): non-magnetic convection). While the resulting hydrodynamic state is linearly stable to magnetic perturbations, it is (nonlinearly) unstable to some finite-amplitude magnetic perturbations that trigger dynamo action. Christensen et al., (2001) suggest that the basin of attraction of this dynamo solution may be small and therefore recommend prescribing a mainly dipolar, initial magnetic field of the form
@@
with (dimensionless) magnetic energy $M_{0}=1215$, which kickstarts an $m=4$ dynamo solution. We note that, as at the benchmark parameters the system is hydrodynamically unstable to multiple modes of convection, other dynamo solutions are possible using different thermal forcing wavenumbers — see Feudel et al., (2017) for instance. In fact, we have checked that our procedure works without modification for identifying solutions with $m=5$ symmetries by replacing the $Y_{4}^{4}$ spherical harmonic in (27) with $Y_{5}^{5}$. However, as the overall dynamical picture is identical to the $m=4$ case, we do not present it here.
Label K M $\omega$ Stable Growth rate
TW0 $58.35$ 00 $0.000182$ Yes -2.06 (-0.80)
TW1 $30.773$ $626.4$ $-0.00310$ Yes -0.56
TW2 $29.72$ $459.9$ $-0.00250$ No 1.13
Table 1: Travelling wave solutions to the governing equations with $m=4$ symmetry. $\omega$ is the frequency at which they drift relative to the reference frame. $K$ and $M$ give their dimensionless kinetic and magnetic energies, respectively. Also displayed are the (dimensionless) least stable growth rates of the solutions to perturbations. For TW0, the growth rate is displayed for a purely hydrodynamic perturbation, with the growth rate for a purely magnetic perturbation also shown in brackets. With our definition of $\omega$, the numerical values differ from the drift frequencies of Christensen et al., (2001) by a factor of Ek.
@@
Figure 2: Slices of the $m=4$ nonlinear dynamo state (TW1). The dashed line on the equatorial slice indicates where the meridional slices were taken.
@@
Using the Newton-hookstep procedure outlined in §2.3, we find the system to possess three $m=4$ travelling wave (TW) solutions at our working parameters, characterised by the energies and drift frequency (relative to the already-rotating reference frame) displayed in table 1. The table shows one purely hydrodynamic solution (TW0) which corresponds to the ‘case 0’ benchmark of Christensen et al., (2001). This hydrodynamic solution arises via a supercritical Hopf bifurcation at $\tilde{\textit{Ra}}_{c}=55.9$ (see the weakly nonlinear analysis of Skene and Tobias, (2024), for instance), and is linearly stable to both hydrodynamic and magnetic perturbations at our working parameters (with growth rates displayed in table 1). The two other TW solutions are MHD solutions that form at a saddle node bifurcation at $\tilde{\textit{Ra}}\sim 98$ (Feudel et al.,, 2017) connected to the purely convecting solution: one stable (TW1) and corresponding to the dynamo (‘case 1’ of Christensen et al., (2001)); the other unstable (TW2). Figure 2 shows slices of TW1, showing that the nonlinear state is mainly dipolar, with a quadrupolar toroidal part (evident from the radial component of the current). We note here that TW2 has a very similar structure to TW1, which is to be expected since we are close to the saddle node bifurcation, although it shows slightly more spatial localisation than TW1. This localisation becomes more apparent at higher Rayleigh numbers, where the two states start to deviate in structure.
@@
We now let the system evolve to the saturated, purely hydrodynamic, $m=4$ convective solution (TW0). This state is then seeded with a magnetic field of the form given by Christensen et al., (2001) (28)-(29); however, the field amplitude is rescaled to have a given initial magnetic energy $M_{0}$, which is gradually decreased (through a series of repeated simulations) down to the point where the initial field fails to trigger a dynamo. The energy timeseries for two slightly different energy budgets close to the threshold are shown in figure 3. Initially, the evolution is very similar for both cases and is characterised by a transient amplification of the magnetic field, which extracts energy from the convection. This transient growth occurs on a dynamical time scale and is followed by a plateau until around $t\approx 2.5$ (i.e. half an ohmic timescale). At this point the evolutions differ, with the $M_{0}=345$ case achieving further growth in magnetic energy to reach a new stable MHD state (corresponding to the TW1 dynamo) over a few ohmic timescales, and the $M_{0}=344$ case magnetic field decaying back to the purely hydrodynamic state TW0 (for future reference we label the rescaled benchmark initial condition with $M_{0}=344$ RB). Remarkably, the flow state around $t\approx 2.5$ corresponds to the unstable MHD state TW2. By performing a stability calculation of TW2, we confirm that it has a single unstable direction (see table 1). Thus, TW2 is found to be a saddle point with both solutions initially approaching TW2 along its stable directions. The flow can then either approach TW0 or TW1 along this unstable direction, with which travelling wave the flow approaches depending on which side of the saddle the state lies.
@@ 3.2 Optimised seeds
With the baseline behaviour for our system established, we now seek to identify the spatial structure of the lowest-energy magnetic perturbation required to reach the dynamo solution, using the optimisation procedure outlined in §$\S$2.2. In other words, what is the smallest magnetic field initial condition, with the hydrodynamic initial condition fixed as TW0, that leads to TW1? Following Mannix et al., (2022), we initially consider to that end the cost functional with cost integrand
$\mathcal{J}_{I}=\frac{1}{2\textit{Ek}\textit{Pm}}\bm{B}\cdot\bm{B},$ (30)
@@
which corresponds to maximising the time-integrated magnetic energy. As our expected magnetic field is large scale, we initialise the optimisation procedure with a random guess with spherical harmonic degrees up to $\ell=m=2$, and a random radial dependence with Jacobi polynomial degree up to four. We note here, that we have tested random initial guesses up to $\ell=m=4$. The behaviour of the optimisation is initially similar, with the seed magnetic field converging to a large scale solution. However, the less sparse spectra of the $\ell=m=4$ guess causes slower convergence, with the optimisation spending a long time removing the increased amount of fine scale components which build up quicker when using this guess. Even though this guess is still large-scale, the initial strongly nonlinear transient period, in which the magnetic field takes energy from the convection, leads to updates which initially increase the higher-scale components of $\bm{B}_{0}$. We further describe the convergence behaviour of the optimisation routine in appendix C. Based on the quick evolution of transients in the previous section, we start by considering a short optimisation window of $t_{\textit{opt}}=0.2$. The initial magnetic energy budget is specified to be $M_{0}=344$ as this is below the threshold at which a dynamo can be reached from the baseline solution.
@@
Figure 4: Comparison of the magnetic energy (top) and kinetic energy (bottom) evolution with four different initial conditions with $M_{0}=344$. The initial conditions are the optimised seeds obtained by optimising the total magnetic energy with a random initial guess (solid blue line), optimising the total magnetic energy with an initial guess of the rescaled benchmark initial condition (RB) (dotted red line), and optimised seed obtained by optimising the energy in the $m=0$ part of the magnetic field starting from a random guess (dashed-dotted green line). We also show the timeseries obtained without optimisation, starting directly from the rescaled benchmark initial condition RB (dashed orange line). The plots on the left show the energy evolution in the optimisation window $t_{\textrm{opt}}=0.2$, and the plots on the right show the long-time evolution.
@@
Figure 5: Top: Slices of an example optimal seed (non-dynamo) obtained with the total magnetic energy cost functional, with $t_{\textrm{opt}}=0.2$ and $\textit{M}_{0}=344$. Middle: Slices of the optimal seed (dynamo) obtained with the axisymmetric magnetic energy cost functional with $t_{\textrm{opt}}=0.2$ and $\textit{M}_{0}=344$. Bottom: Slices of the optimal seed (dynamo) obtained with the axisymmetric magnetic energy cost functional with $t_{\textrm{opt}}=0.4$ and $\textit{M}_{0}=162$. Figures (a), (d) and (g) show meridional slices the radial component of the magnetic field, figures (b), (e) and (h) show meridional slices of the radial component of the current, and figures (c), (f) and (i) show equatorial slices of the latitudinal component of the magnetic field. The dashed lines on the equatorial slices indicates where the meridional slices were taken.
@@
Increasing the time horizon of the optimisation procedure up to $t_{\textrm{opt}}=0.4$ was not found to add robustness to the computation of optimal seeds as multiple local extrema characterised by strong transient growth (but not necessarily a long-time dynamo) continue to coexist at such target times. While it is tempting to further increase $t_{\textrm{opt}}$ — indeed we expect that optimising over (very) long times would ultimately discriminate between dynamo seeds and magnetic perturbations yielding only transient growth and ultimately decay -, for long target times convergence becomes increasingly difficult (numerical cost notwithstanding). It is therefore more realistic to define another cost functional that could possibly reduce the number of local extrema, yielding more robust seed identification. This is consistent with the observation of Kerswell et al., (2014) that “the measure used as the objective functional and the norm constraining the initial condition do not need to be the same" for computing minimal seeds.
@@
$\mathcal{J}_{I}=\frac{1}{2\textit{Ek}\textit{Pm}}\bar{\bm{B}}\cdot\bar{\bm{B}},$ (31)
@@
The result of optimising this cost functional is also shown in figure 4 (dashed-dotted, green line), and we label this optimal seed OS1. In contrast to optimising the total magnetic energy, optimising this new cost functional is robust in the sense that it was found to always identify the same optimal seed (regardless of the random initial guess used to initialise the algorithm). Furthermore, the computed optimal seed does indeed lead to a dynamo. Incidentally, it is also found to outperform some “optimal” seeds that (locally) maximise the total magnetic energy (as exemplified in figure 4). Based on these observations, one could perhaps speculate that using this new cost functional might be a way to ‘smooth out’ the optimisation landscape, possibly suppressing the multiple local extrema and thus providing a good proxy for the global extremum of the first (total energy) optimisation problem — although this is highly uncertain.
@@
Figure 6: Robustness of the optimisation results with respect to the time horizon $t_{\textrm{opt}}$. Left: The minimum initial magnetic energy budget to find a dynamo solution for a given time horizon $t_{\textrm{opt}}$ ($m=0$ cost functional). Circles show runs that succeed in finding a dynamo, and crosses indicate no dynamo was found. Right: Magnetic energy timeseries for $M_{0}=165$ starting from optimised seeds ($m=0$ cost functional) computed with different time horizons.
@@
With a robust cost functional identified, we now turn our attention to lowering the energy of the initial seed. To this end, we perform a series of optimisations maximising the total axisymmetric energy, for gradually decreasing initial energy budgets at fixed time horizon. Each optimal seed then serves as initial condition for a long-time DNS: if a dynamo state is found at long times, then $M_{0}$ is decreased further and a new optimisation is performed. We thus determine the initial energy $M_{0}^{*}$ below which the computed optimal seed no longer triggers a subcritical dynamo, and repeat the whole procedure for increasing time horizon. Indeed Pringle and Kerswell, (2010); Kerswell et al., (2014) already noted that the choice of time horizon is crucial in accurately determining the amplitude of the minimal seed for transition to turbulence. As shown in Figure 6 (Left), $M_{0}^{*}$ was found not to vary much as the target time was increased up to $t_{\textrm{opt}}=0.4$. Moreover, the effect of lengthening the optimisation time horizon is shown in Figure 6 (Right) for an initial energy budget of $\textit{M}_{0}=165$: at this energy $t_{\textrm{opt}}=0.2$ is not long enough to optimise an initial condition which yields a dynamo. We see that lengthening the time horizon to $t_{\textrm{opt}}=0.3$ and $t_{\textrm{opt}}=0.4$ enables the optimisation procedure to converge to a dynamo solution. Although the early timeseries in Figure 6 (Right) show that there are differences in the magnetic energy evolution as the $t_{\textrm{opt}}$ is changed, we see that these changes become significantly less as $t_{\textrm{opt}}$ is increased. This indicates that the computed extremum eventually becomes insensitive to $t_{\textrm{opt}}$, which we can attribute to the fact that for $t_{\textrm{opt}}\approx 0.4$ the transient growth regime has ended and the flow has approached the unstable MHD state TW2. For longer time-horizons the optimisation procedure struggles to converge.
@@
Based on these findings we conclude that $t_{\textrm{opt}}=0.4$ is the most practical choice of time horizon, enabling the optimisation procedure to converge, while at the same time being long enough to ensure robustness with respect to changes in $t_{\textrm{opt}}$. Recall, this is non-dimensionalised with respect to the viscous diffusion time, and this is a dynamical timescale. In other words, we only need to optimise up to when the state approaches the unstable edge state, at which point we require the instability of the edge state to lead to the dynamo solution. This is similar to previous studies finding minimal seeds for flows with edge states (see the work of Duguet et al., (2010); Juniper, (2011) for example). The spatial structure of the ‘minimal’ seed (i.e. the lowest-energy dynamo seed we could find) is shown in figure 5 (Bottom) ; it is similar to the optimal seed found at higher energy (OS1) $\textit{M}_{0}=344$ (figure 5, Middle), however shows more localisation (indicating a more concentrated magnetic energy density profile in all spatial directions). We note that neither of these magnetic fields closely resemble either the nonlinear dynamo state (figure 2) or unstable state (similar to figure 2). We see that although they have some similar features, such as a dominant dipolar poloidal part, and quadrupolar toroidal part, the radial structure and energy spectra is significantly different between the two solutions - and is more complex in the case of the optimal seed. This showcases the need for a systematic procedure that can work from random initial guesses (Mannix et al.,, 2022). The minimal seed, obtained with $M_{0}=162$ and $t_{\textrm{opt}}=0.4$ is labelled OS2. Additionally, we also label the optimal seed obtained with $M_{0}=161$ and $t_{\textrm{opt}}=0.4$, which does not reach a dynamo solution, OS3.